if the perimeter of a rectangle is 108 inches, and its lenght is 3 1/4 feet, find it width.
Answer by tutorcecilia(2144)

Use the formula for the perimeter of a rectangle:
P=2(length + width)
Plug-in the values
108=2(3 1/4 + width)
Change the fraction to a decimal
108=2(3.25+width) [solve for w]
108/2=3(3.25+w)/2
54=3.25+w
54-3.25+2
50.75=w
.
check by plugging all of the values back into the original equation:
P=2(length + width)
108=2(3.25+50.75
108=2(54)
108=108 [checks out]